A pair of dice are rolled​ 1,000 times with the following frequencies of​ outcomes:

Sum

2

3

4

5

6

7

8

9

10

11

12

Frequency

10

30

50

70

110

150

170

140

120

80

70

Use these frequencies to calculate the approximate empirical probabilities and odds for the events a. The sum is less than

5

or greater than

8.

b. The sum is even or exactly divisible by

3.
Answer #1

(a) sum is less than 5 or greater than 8 sum less than 5 is P(g) + P(3) + P(4) = 10 + 30+50 = 90 1000 1000 sum greater than 8(o sum is even on exactly divisible by 3. d DLA = Sum is even =) p (2) + P(4)+ P(6)+ P(8)+ P(10) + P(92) 10+50+11ot 170+ 120+sum TO is even and exactly divisible by 3, au 120.26 P(12) = 10 + 110 1000 1000 Sum is even of exactly divisible by 3 y 930 +